4. 2 points PSE6 4.P049 My Notes A ball on the end of a string is whirled around in a horizontal circle of radius 0.340 m. The plane of the circle is 1.00 m above the ground. The string breaks and the ball lands 1.60 m (horizontally) away from the point on the ground directly beneath the balls location when the string breaks. Find the radial acceleration of the ball during its circular motion. Magnitude m/s Direction away from the center of curvature toward the center of curvatureIf the number is a decimal please include 4 digits past the decimal point if available.
Answer #1

Time taken between breaking the string and landing on ground =t=(2h/g).5=0.45s

g=acceleration due to gravity=9.8 m/s2

Horizontal distance=x=vt=1.6

V=3.56 m/s

So magnitude of radial acceleration=a=v2/r=36.8941m/s2